Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.)
cos 2θ = cos^2 θ − 1/2


θ =

Answer by lwsshak3(11628)

Solve the given equation. (Enter your answers as a comma-separated list. Let k be any integer. Round terms to three decimal places where appropriate. If there is no solution, enter NO SOLUTION.)
cos 2θ = cos^2 θ − 1/2
cos^2x-sin^2x=cos^2x-1/2
sin^2x=1/2
sinx=1/√2=√2/2
x=π/4+2πk, 3π/4+2πk, k=integer